# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024ccs: Computing methodologies Natural language processingccs: Applied computing Computational biology
@@ 1. Introduction
To facilitate the understanding of scientific languages, researchers have devised Scientific Large Language Models (Sci-LLMs) customized for various scientific domains and disciplines. For instance, molecular language models have been developed to represent molecule structures as a string of atoms and chemical bonds (Li and Jiang, 2021). These models aid in predicting molecular properties (Wang et al., 2019), designing new drugs (Zhang et al., 2022b), and proposing retrosynthesis routes (Schwaller et al., 2021). Similarly, protein language models operate based on sequences of amino acids (Rives et al., 2019; Brandes et al., 2022). They are used to forecast 3D protein structures and functions (Lin et al., 2022), enhance existing proteins for improved fitness (Notin et al., 2023b), and create new proteins with specific functionalities (Nijkamp et al., 2023). As a burgeoning area within the AI-for-Science research, many Sci-LLMs have been proposed with modified architectures, learning methods, training corpus, and evaluation benchmarks and criteria. Despite their notable achievements, these models have mostly been explored within their respective research domains. There is currently a lack of a comprehensive review that unifies these language modeling advancements.
@@
This survey’s distinct boundaries set it apart from other reviews of LLMs and computational modeling for molecules, proteins, and genomes. In contrast to those primarily centered on natural languages (Zhao et al., 2023; Yang et al., 2023), our emphasis leans more towards scientific languages. Unlike surveys concentrating solely on molecule (Du et al., 2022; Xia et al., 2023), protein (Bepler and Berger, 2021; Hu et al., 2022; Tran et al., 2023; Unsal et al., 2022), or genome data (Consens et al., 2023), we aim to provide a comprehensive view of language models for chemical and biological research. Furthermore, we delve into multimodal LLMs, exploring the interactions between texts and molecule/protein/genome languages. To the best of our knowledge, this nuanced exploration has not been covered in previous surveys. The contributions of the survey can be summarized as follows:
@@
Figure 3. An evolutionary tree of Sci-LLMs, which consists of five main branches corresponding to the research scopes in this survey. Due to the extensive number of Sci-LLMs, it is not feasible to include all of them in this figure, despite their exceptional quality. For detailed information on the featured models, please refer to Table 1, 3, 5, 7, and 9. Additionally, we will share the source file of this figure and encourage readers to make incremental updates at https://kdocs.cn/l/cbRA94QwMhmn.
@@ 3. Textual Scientific Large Language Models
(Sec. 3.1),text width=4em [Biological, text width=4em [Encoder-only: BioBERT (Lee et al., 2020)   BlueBERT (Peng et al., 2020)  PubMedBERT (Gu et al., 2021)  BioMegatron (Shin et al., 2020) ] [Decoder-only: BioGPT (Luo et al., 2022b)   BioMedGPT-LM (Luo et al., 2023c) ] ] [Chemical, text width=4em [Encoder-only: ChemBERT (Guo et al., 2021) MatSciBERT (Gupta et al., 2022) MaterialsBERT (Shetty et al., 2023) ] ] [Comprehen., text width=4em [ Encoder-only: SciBERT (Beltagy et al., 2019) ScholarBERT (Hong et al., 2023) ] [Decoder-only: DARWIN-Base  (Xie et al., 2023) ] ] ] [Datasets
(Sec. 3.2),text width=4em [Pre-training, text width=4em [ PubMed   PubChem   bioRxiv   arXiv   WIPO  etc. ] ] [Benchmarks, text width=4em [PubMedQA (Jin et al., 2019)   MMLU (Hendrycks et al., 2021)   C-Eval (Huang et al., 2023)   AGIEval (Zhong et al., 2023) ScienceQA (Lu et al., 2022)
Bioinfo-Bench-QA (Chen and Deng, 2023)   SciQ (Johannes Welbl, 2017)  Xiezhi (Gu et al., 2023)   ARC (Clark et al., 2018)   SciEval (Sun et al., 2023) ] ] ] [Evaluation
@@ 3.1. Models
To provide a clear overview of Text-Sci-LLMs, we divide them into three distinct categories: biological, chemical, and comprehensive, based on the specific scientific domains they focus on. Table 1 shows the summary of Text-Sci-LLMs.
@@
Biological Domain. Large language models trained on extensive biological corpora, such as BERT (Devlin et al., 2019) and its variations with the encoder-only LLM architecture, have demonstrated significant potential in Natural Language Processing (NLP) tasks within biology. BioBert (Lee et al., 2019), BioMegatron (Shin et al., 2020), PubMedBERT (Gu et al., 2021), BioM-BERT (Alrowili and Shanker, 2021) and LinkBERT (Yasunaga et al., 2022) initially trained on broad corpora like Wikipedia and textbooks and then fine-tuned on specific biological NLP tasks, showcase their significant improvements in various downstream tasks such as biological terminology understanding, named entity recognition, text similarity, and relation extraction. BlueBERT (Peng et al., 2020) presented a thorough empirical study on the efficacy of multi-task learning on BERT-based models for biomedical texts. Since the encoder-only models lack the capability of textual information generation, GPT and its variants (Radford et al., 2018, 2019a; Brown et al., 2020) with a decoder-only architecture have emerged as dominant players in the field of biological NLP. BioGPT (Luo et al., 2022b), an extension of GPT-2 (Radford et al., 2019b), has been extensively pretrained on biomedical literature, showcasing remarkable performance in relation extraction and question answering. Moreover, it demonstrates the ability to generate coherent and fluent descriptions within the biomedical context. BioMedGPT-LM (Luo et al., 2023c), which is incrementally pretrained on LLaMA2 (Touvron et al., 2023a), enabling the comprehensive understanding of various biological modalities and aligning them with natural language.
@@
Comprehensive Domain. This part introduces general Text-Sci-LLMs, which span diverse domains including biology, chemistry, and other scientific disciplines. SciBERT (Beltagy et al., 2019) is built upon the BERT architecture and utilizes unsupervised pretraining on a vast corpus of scientific publications to improve performance on downstream scientific NLP tasks. ScholarBERT (Hong et al., 2023) is a series of BERT-based scientific language models trained on a substantial and diverse scientific corpus spanning domains including life sciences, biomedicine and physical sciences. DARWIN-Base (Xie et al., 2023), based on a decoder-only architecture, is trained on a massive corpus from various scientific texts and knowledge bases. Notably, it was fine-tuned with instruction data from the SciQ (Welbl et al., 2017) dataset and additional question-answer pairs, exhibiting proficiency in tasks that require extensive scientific knowledge. We emphasize that, although the general LLMs such as ChatGPT (Ouyang et al., 2022), GPT4 (OpenAI, 2023) and Gemini (Team, 2023) encompass a wealth of scientific knowledge, they are not tailored for Sci-LLMs, and therefore do not fall under the category of general Text-Sci-LLMs.
@@ 3.2. Datasets
  • PubMedQA (Jin et al., 2019). PubMedQA is a biomedical Question-Answering (QA) dataset collected from PubMed, which has 1k expert-annotated, 61.2k unlabeled and 211.3k artificially generated QA instances.
@@
Dataset Time Subset #Item Domain Type Capability Language
@@
Basic-hemistry 2909 Chemistry
@@
Bioinfo-Bench-QA (Chen and Deng, 2023) 2023.10 - 150 Biology Multiple choice Post-college English
BLURB (Gu et al., 2021) 2020.07 - 648k Biology Multiple NLP tasks Mixed
PubMedQA (Jin et al., 2019) 2019.09 - 273.2k Biology True or false College

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024
@@ 1. Introduction
To facilitate the understanding of scientific languages, researchers have devised Scientific Large Language Models (Sci-LLMs) customized for various scientific domains and disciplines. For instance, molecular language models have been developed to represent molecule structures as a string of atoms and chemical bonds (Li and Jiang, 2021). These models aid in predicting molecular properties (Wang et al., 2019), designing new drugs (Zhang et al., 2022c), and proposing retrosynthesis routes (Schwaller et al., 2021). Similarly, protein language models operate based on sequences of amino acids (Rives et al., 2019; Brandes et al., 2022). They are used to forecast 3D protein structures and functions (Lin et al., 2022), enhance existing proteins for improved fitness (Notin et al., 2023b), and create new proteins with specific functionalities (Nijkamp et al., 2023). As a burgeoning area within the AI-for-Science research, many Sci-LLMs have been proposed with modified architectures, learning methods, training corpus, and evaluation benchmarks and criteria. Despite their notable achievements, these models have mostly been explored within their respective research domains. There is currently a lack of a comprehensive review that unifies these language modeling advancements.
@@
This survey’s distinct boundaries set it apart from other reviews of LLMs and computational modeling for molecules, proteins, and genomes. In contrast to those primarily centered on natural languages (Zhao et al., 2023b; Yang et al., 2023a), our emphasis leans more towards scientific languages. Unlike surveys concentrating solely on molecule (Du et al., 2022; Xia et al., 2023), protein (Bepler and Berger, 2021; Hu et al., 2022; Tran et al., 2023; Unsal et al., 2022), or genome data (Consens et al., 2023), we aim to provide a comprehensive view of language models for chemical and biological research. Furthermore, we delve into multimodal LLMs, exploring the interactions between texts and molecule/protein/genome languages. To the best of our knowledge, this nuanced exploration has not been covered in previous surveys. The contributions of the survey can be summarized as follows:
@@
Figure 3. An evolutionary tree of Sci-LLMs, which consists of five main branches corresponding to the research scopes in this survey. Due to the extensive number of Sci-LLMs, it is not feasible to include all of them in this figure, despite their exceptional quality. For detailed information on the featured models, please refer to Table 1, 3, 5, 7, and 9. Considering the rapid development of Sci-LLMs, we will share the source file of this figure and encourage readers to make incremental updates at https://kdocs.cn/l/cbRA94QwMhmn.
@@ 3. Textual Scientific Large Language Models
(Sec. 3.1),text width=4em [Medical, text width=4em [Encoder-only: ClinicalBERT (Huang et al., 2020)  GatorTron (Yang et al., 2022a)  ] [Decoder-only: BioMedLM (Venigalla et al., 2022)  GatorTronGPT (Peng et al., 2023)  MEDITRON (Chen et al., 2023a)  Meerkat-7B (Kim et al., 2024)
ClinicalGPT (Wang et al., 2023f)  Qilin-Med (Ye et al., 2023)  ChatDoctor (Li et al., 2023d)  BenTsao (Wang et al., 2023d)  HuatuoGPT (Zhang et al., 2023)
Baize-healthcare (Xu et al., 2023a)  zhongjing (Yang et al., 2023b)  PMC-LLaMA (Wu et al., 2023a)  CPLLM (Shoham and Rappoport, 2023)  Clinical Camel (Toma et al., 2023)
Med-PaLM 2 (Singhal et al., 2023b)  Me-LLaMA(Xie et al., 2024)  BiMediX(Pieri et al., 2024)  ] [Encoder-decoder: DoctorGLM (Xiong et al., 2023)  BianQue (Chen et al., 2023d)  Medical-mT5(García-Ferrero et al., 2024) ] ] [Biological, text width=4em [Encoder-only: BioBERT (Lee et al., 2020)   BlueBERT (Peng et al., 2020)  PubMedBERT (Gu et al., 2021)  BioMegatron (Shin et al., 2020) BioLinkBERT (Yasunaga et al., 2022) ] [Decoder-only: BioGPT (Luo et al., 2022b)   BioMedGPT-LM (Luo et al., 2023c) BioELMo (Sharma et al., 2019) BioinspiredLLM (Luu and Buehler, 2023) BioMistral (Labrak et al., 2024) BioMed-RoBERTa (Gururangan et al., 2020) ] ] [Chemical, text width=4em [Encoder-only: ChemBERT (Guo et al., 2021) MatSciBERT (Gupta et al., 2022) MaterialsBERT (Shetty et al., 2023) PharmGPT (Chen et al., 2024b) ] ] [Comprehen., text width=4em [ Encoder-only: SciBERT (Beltagy et al., 2019) ScholarBERT (Hong et al., 2023) INDUS (Bhattacharjee et al., 2024) ] [Decoder-only: DARWIN-Base  (Xie et al., 2023) ] [Encoder-decoder: SciGLM  (Zhang et al., 2024a) ] ] ] [Datasets
(Sec. 3.2),text width=4em [Pre-training, text width=4em [ PubMed   PubChem   bioRxiv   arXiv   WIPO  etc. ] ] [Benchmarks, text width=4em [PubMedQA (Jin et al., 2019b)   MMLU (Hendrycks et al., 2021)   C-Eval (Huang et al., 2023)   AGIEval (Zhong et al., 2023) ScienceQA (Lu et al., 2022)
Bioinfo-Bench-QA (Chen and Deng, 2023)   SciQ (Johannes Welbl, 2017)  Xiezhi (Gu et al., 2023)   ARC (Clark et al., 2018)   SciEval (Sun et al., 2023) SciKnowEval (Feng et al., 2024) GAOKAO-Bench (Zhang et al., 2023c) SciAssess (Cai et al., 2024a) ] ] ] [Evaluation
@@ 3.1. Models
To provide a clear overview of Text-Sci-LLMs, we divide them into four distinct categories: medical, biological, chemical, and comprehensive, based on the specific scientific domains they focus on. Table 1 shows the summary of Text-Sci-LLMs.
Medical Domain. Clinical medicine profoundly influences human health by diagnosing, treating, and preventing diseases. Recent years have witnessed remarkable advances in artificial intelligence technology, leading to the emergence of specialized large language models such as BioMedLM(Venigalla et al., 2022), GatorTron(Yang et al., 2022a), GatorTronGPT(Peng et al., 2023), PMC-LLaMA(Wu et al., 2023a), MEDITRON(Chen et al., 2023a) and Meerkat-7B(Kim et al., 2024). These models are trained with professional materials such as medical literature, clinical records, and medical textbooks, thus significantly advancing the medical industry.
Models like Med-PaLM 2(Singhal et al., 2023b), HuatuoGPT(Zhang et al., 2023), BianQue(Chen et al., 2023d), ClinicalGPT(Wang et al., 2023f), DoctorGLM(Xiong et al., 2023), chatDoctor(Li et al., 2023d), Baize-healthcare(Xu et al., 2023a), zhongjing(Yang et al., 2023b), BenTsao(Wang et al., 2023d), Clinical Camel(Toma et al., 2023), and the comprehensive Me-LLaMA(Xie et al., 2024) have facilitated more efficient communication between doctors and patients through training and fine-tuning in real-world medical records and doctor-patient dialogue datasets. Moreover, the landscape of medical AI has expanded with the introduction of multi-lingual models like BiMediX(Pieri et al., 2024) and Medical-mT5(García-Ferrero et al., 2024), which facilitate medical interactions across different languages, enhancing the accessibility of medical AI globally.
The innovative Qilin-Med(Ye et al., 2023) stands out with its three-stage training method, addressing the need for specialized knowledge in the medical field and offering insights into tailored training for professional domains. Additionally, models like ClinicalBERT(Huang et al., 2020) predict 30-day readmissions, while CPLLM(Shoham and Rappoport, 2023) forecasts future disease risks using patients’ historical diagnoses, offering doctors scientific references for diagnosis and treatment decisions. Continuous advances in these technologies bring renewed optimism to the field of medicine and make substantial contributions to advancing human health endeavors.
@@
Biological Domain. Large language models trained on extensive biological corpora, such as BERT (Devlin et al., 2019) and its variations with the encoder-only LLM architecture, have demonstrated significant potential in Natural Language Processing (NLP) tasks within biology. BioBert (Lee et al., 2020), BioMegatron (Shin et al., 2020), PubMedBERT (Gu et al., 2021), BioM-BERT (Alrowili and Shanker, 2021), BioELMo (Sharma et al., 2019) and LinkBERT (Yasunaga et al., 2022) initially trained on broad corpora like Wikipedia and textbooks and then fine-tuned on specific biological NLP tasks, showcase their significant improvements in various downstream tasks such as biological terminology understanding, named entity recognition, text similarity, and relation extraction. BlueBERT (Peng et al., 2020) presented a thorough empirical study on the efficacy of multi-task learning on BERT-based models for biomedical texts. Since the encoder-only models lack the capability of textual information generation, GPT and its variants (Radford et al., 2018, 2019a; Brown et al., 2020) with a decoder-only architecture have emerged as dominant players in the field of biological NLP. BioGPT (Luo et al., 2022b), an extension of GPT-2 (Radford et al., 2019b), has been extensively pretrained on biomedical literature, showcasing remarkable performance in relation extraction and question answering. Moreover, it demonstrates the ability to generate coherent and fluent descriptions within the biomedical context. BioMedGPT-LM (Luo et al., 2023c), which is incrementally pretrained on LLaMA2 (Touvron et al., 2023a), enabling the comprehensive understanding of various biological modalities and aligning them with natural language. BioinspiredLLM (Luu and Buehler, 2023), based on the Llama2 model and enhanced with the Orca-2-13b model (Mitra et al., 2023), specializes in elucidating the mechanics of biological and bio-inspired materials, exhibiting cognitive capabilities such as information retrieval, hypothesis generation, and collaborative interactions with other AI models. BioMistral (Labrak et al., 2024) is based on the Mistral model and further pre-trained on the corpus sourced from PubMed Central, thereby enhancing its performance in medical question-answering tasks. BioMed-RoBERTa (Gururangan et al., 2020) is a language model based on the RoBERTa-base architecture. It is adapted through continued pretraining on 2.68 million scientific full-text papers from the Semantic Scholar corpus.
@@
Medical ClinicalBERT (Huang et al., 2020) 2019.04 110M BERT MIMIC-III
GatorTron (Yang et al., 2022a) 2022.02 8.9B BERT MIMIC-III, PubMed, etc.
BioMedLM (Venigalla et al., 2022) 2022.12 2.7B GPT PubMed
GatorTronGPT (Peng et al., 2023) 2023.05 5B, 20B GPT Pubmed, Custom Data
MEDITRON (Chen et al., 2023a) 2023.11 7B, 70B LLAMA2 Pubmed, PMC, etc.
Meerkat-7B (Kim et al., 2024) 2024.03 7B Mistral Custom Data $\times$
ClinicalGPT (Wang et al., 2023f) 2023.06 7B BLOOM cMedQA2, MedDialog, etc. $\times$
Qilin-Med (Ye et al., 2023) 2023.10 7B Baichuan ChiMed
ChatDoctor (Li et al., 2023d) 2023.03 7B LLAMA HealthCareMagic
BenTsao (Wang et al., 2023d) 2023.04 7B LLAMA Custom Data
HuatuoGPT (Zhang et al., 2023) 2023.05 7B BLOOM Custom Data
Baize-healthcare (Xu et al., 2023a) 2023.04 7B LLAMA MedQuAD
zhongjing (Yang et al., 2023b) 2023.08 13B LLAMA Medical Books, Wiki, etc.
PMC-LLaMA (Wu et al., 2023a) 2023.04 13B LLAMA PMC, Medical books, etc.
CPLLM (Shoham and Rappoport, 2023) 2023.09 2.7B, 13B LLAMA-2 eICU-CRD, MIMIC-IV
Clinical Camel (Toma et al., 2023) 2023.05 13B, 70B LLAMA-2 PubMed $\times$
Med-PaLM 2 (Singhal et al., 2023b) 2023.05 340B PaLM2 MultiMedQA $\times$
DoctorGLM (Xiong et al., 2023) 2023.04 6.2B ChatGLM CMD., HealthCareMagic, etc.
BianQue (Chen et al., 2023d) 2023.10 6.2B ChatGLM BianQueCorpus
Medical-mT5 (García-Ferrero et al., 2024) 2024.01 738M, 3B T5 Custom Data $\times$
Me-LLaMA (Xie et al., 2024) 2024.02 13B, 70B LLAMA Custom Data
BiMediX (Pieri et al., 2024) 2024.02 7B Mixtral BiMed1.3M
BioELMo (Jin et al., 2019a) 2019.04 - ELMo PubMed
@@
BioinspiredLLM (Luu and Buehler, 2023) 2023.09 13B Llama-2 Biological article
BioMistral (Labrak et al., 2024) 2024.02 7B Mistral PMC
@@
PharmGPT (Chen et al., 2024b) 2024.02 13B, 70B LLaMA Paper, report, book, etc. $\times$
@@
SciGLM (Zhang et al., 2024a) 2024.01 6B, 32B ChatGLM3 SciInstruct
Uni-SMART (Cai et al., 2024b) 2024.03 7B - Patents, news, literature, etc. $\times$
INDUS (Bhattacharjee et al., 2024) 2024.05 125M RoBERTa wikipedia, PubMed, PMC, etc. $\times$
@@
Comprehensive Domain. This part introduces general Text-Sci-LLMs, which span diverse domains including biology, chemistry, and other scientific disciplines. SciBERT (Beltagy et al., 2019) is built upon the BERT architecture and utilizes unsupervised pretraining on a vast corpus of scientific publications to improve performance on downstream scientific NLP tasks. ScholarBERT (Hong et al., 2023) is a series of BERT-based scientific language models trained on a substantial and diverse scientific corpus spanning domains including life sciences, biomedicine and physical sciences. INDUS (Bhattacharjee et al., 2024) is a suite of language models tailored for scientific applications in domains such as Earth science, biology, physics, heliophysics, planetary sciences, and astrophysics, trained using curated scientific corpora. It includes encoder models for natural language understanding tasks and contrastive-learning-based text embedding models for information retrieval, with smaller, efficient versions created using knowledge distillation techniques. DARWIN-Base (Xie et al., 2023), based on a decoder-only architecture, is trained on a massive corpus from various scientific texts and knowledge bases. Notably, it was fine-tuned with instruction data from the SciQ (Welbl et al., 2017) dataset and additional question-answer pairs, exhibiting proficiency in tasks that require extensive scientific knowledge. We emphasize that, although the general LLMs such as ChatGPT (Ouyang et al., 2022), GPT-4 (OpenAI, 2023) and Gemini (Team, 2023) encompass a wealth of scientific knowledge, they are not tailored for Sci-LLMs, and therefore do not fall under the category of general Text-Sci-LLMs. SciGLM (Zhang et al., 2024a) is a suite of scientific language models fine-tuned from the ChatGLM family, using a novel self-reflective instruction annotation framework. This innovative method generates high-quality reasoning data, enabling SciGLM to excel in scientific reasoning and numerical calculations. Uni-SMART (Cai et al., 2024b) is a model designed to interpret and analyze multimodal scientific literature, including text, tables, charts, and molecular structures. It uses key methods such as multimodal learning, iterative training, and self-reflective instruction annotation to enhance performance in scientific tasks.
@@ 3.2. Datasets
  • The Medical Information Mart for Intensive Care (MIMIC) series, including MIMIC-III (Johnson et al., 2016), MIMIC-IV (Johnson et al., 2023a), MIMIC-CXR (Johnson et al., 2019), and MIMIC-IV-Note (Johnson et al., 2023b), are publicly available medical databases derived from the Beth Israel Deaconess Medical Center.
    • MIMIC-III (Johnson et al., 2016) is a de-identified database containing data from over 40,000 critically ill patients between 2001 and 2012, and it has been extensively utilized in research areas such as epidemiology, machine learning, and clinical informatics.
    • MIMIC-IV (Johnson et al., 2023a) represents an updated version of MIMIC-III (Johnson et al., 2016), offering medical data from 2008 to 2019, further enhancing the depth and breadth of healthcare research.
    • MIMIC-CXR (Johnson et al., 2019) focuses on chest X-ray imaging and includes 377,110 chest X-ray images along with their corresponding radiological text reports. It has leveraged natural language processing tools to extract 14 labels from these reports, providing a significant impetus for the development of medical computer vision technology.
    • MIMIC-IV-Note (Johnson et al., 2023b), contains de-identified discharge summaries for 331,794 patients, totaling 145,915 documents, and radiological reports for 2,321,355 patients, totaling 237,427 documents, offering a rich dataset for clinical research in hospitals and emergency departments.
  • The eICU Collaborative Research Database (eICU-CRD) (Pollard et al., 2018) is a publicly available, multi-center critical care database jointly created by Philips Healthcare and the Laboratory for Computational Physiology at the Massachusetts Institute of Technology (MIT). It contains detailed medical data from over 200,000 ICU patients in the United States.
  • The cMedQA2 (Zhang et al., 2018) dataset is a Chinese medical question-and-answer dataset that is aggregated from a Chinese medical question-and-answer online forum [^10] http://www.xywy.com/, consisting of 120,000 questions and 226,000 answers.
  • MedDialog (He et al., 2020) is a large-scale medical dialogue dataset comprising 1.1 million conversations and 4 million utterances between patients and doctors.
  • The ChiMed (Ye et al., 2023) dataset is a large-scale Chinese dataset designed for training Large Language Models (LLMs) in the medical field. It includes various types of data such as medical question-and-answer, texts, knowledge graphs, and dialogues. The dataset is divided into three stages: Domain-Specific Continuous Pre-training (DCPT), Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO).
  • The HealthCareMagic-100k is a dataset composed of approximately one million real-life doctor-patient conversations, sourced from the online medical consultation website HealthCareMagic111111www.healthcaremagic.com.
  • The MedQuAD (Ben Abacha and Demner-Fushman, 2019) dataset is for medical question-and-answer systems and supports Recognizing Question Entailment (RQE) methods. It has 47,457 medical question-and-answer sets with annotations from trusted NIH-related sources.
  • MultiMedQA (Singhal et al., 2023a) is a comprehensive benchmark dataset for medical question-answering that integrates multiple existing datasets from medical examinations, medical research, and consumer health inquiries, and also includes the newly added HealthSearchQA dataset, which contains 3,173 commonly searched health-related questions from users.
  • The OpenI (Demner-Fushman et al., 2016) dataset collected 8,121 images from the Picture Archiving Systems (PACS) of Indiana University Hospitals, along with 3,996 corresponding radiology reports. To improve the relevance of the retrieved clinical documents, manual coding was also added to the radiologists’ reports.
  • Psych8k (Liu et al., 2023a) is a dataset designed for the field of psychological counseling, based on 260 in-depth individual interviews, each lasting one hour, with an average of 31 question-and-answer pairs per individual. It covers a variety of mental health topics, including anxiety, depression, stress, and interpersonal relationships, and also takes into account minority groups.
  • The CMD. dataset (Xiong et al., 2023) is a multi-departmental Chinese medical dialogue dataset that compiles medical consultation records from various departments including surgery, obstetrics and gynecology, pediatrics, internal medicine, and breast surgery, totaling over 900,000 question-and-answer pairs.
  • BianQueCorpus (Chen et al., 2023d) is a dataset of multi-turn questioning and health advice that has been polished by ChatGPT. It contains over 2.4 million samples, with questions accounting for 46.2% of the doctor’s responses, aimed at helping the model better understand patient conditions and provide targeted suggestions.
  • ChemData(Zhang et al., 2024b) is a large-scale dataset for tuning language models in chemistry, encompassing nine core chemistry tasks and featuring 730,000 high-quality questions and answers, sampled from a subset of 700,000 data pieces.
  • SciInstruct (Zhang et al., 2024a) is a comprehensive scientific instruction tuning dataset. It contains 254k data entries including physics, chemistry problems, math, and formal proofs.
@@
  • MedQA-USMLE (Jin et al., 2020) dataset covers USMLE exam questions collected from the web, comprising a total of 10,178/1,272/1,273 train/dev/test standard questions that describe medical scenarios, which doctors should be able to answer, along with four possible answer choices provided for selection.
  • MedMCQA (Pal et al., 2022) is a large-scale multisubject multiple-choice dataset, comprising over 194,000 high-quality questions from Indian medical entrance exams like AIIMS and NEET. This dataset addresses 2.4k healthcare topics in 21 medical topics.
  • JAMA Clinical Challenge dataset (Chen et al., 2024a). The JAMA Clinical Challenge dataset is a medical question-answer dataset composed of 1524 real and challenging clinical cases, covering 13 medical fields, with each case including a question, four answer choices, and detailed explanations provided by experts.
  • CMtMedQA (Yang et al., 2023b) is a Chinese medical dialogue dataset consisting of 70,000 authentic multiturn doctor-patient conversations across 14 medical departments, designed to enhance the dialogue and proactive inquiry capabilities of large language models in the field of traditional Chinese medicine.
  • huatuo-26M (Li et al., 2023f) is a large-scale Chinese medical question-answering (QA) dataset, containing 26 million QA pairs. The dataset primarily sources from online medical consultation websites, medical encyclopedias, and medical knowledge bases.
@@
  • GAOKAO-Bench (Zhang et al., 2023c) is a benchmark designed to evaluate the performance of large language models (LLMs) using questions from the Chinese GAOKAO examination. It includes both subjective and objective questions and aims to provide a comprehensive and human-aligned assessment of LLM capabilities across various subjects.
@@
  • PubMedQA (Jin et al., 2019b). PubMedQA is a biomedical Question-Answering (QA) dataset collected from PubMed, which has 1k expert-annotated, 61.2k unlabeled and 211.3k artificially generated QA instances.
@@
  • SciBench (Wang et al., 2023c) is an academic benchmark encompassing physics, chemistry, and mathematics. It comprises a comprehensive set of 695 problems derived from instructional textbooks, specifically tailored for college-level scientific problem-solving. The primary objective of this benchmark is to assess complex reasoning capabilities, strong domain knowledge, and advanced calculation skills of LLMs.
  • SciAssess (Cai et al., 2024a) is a benchmark designed to evaluate LLMs’ proficiency in scientific literature analysis across three levels: Memorization, Comprehension, and Analysis & Reasoning, using tasks from diverse scientific fields. It aims to thoroughly assess LLM capabilities and support their development in scientific literature analysis.
  • SciKnowEval (Feng et al., 2024) is a benchmark for evaluating large language models’ scientific knowledge and abilities across five progressive levels, including knowledge coverage, inquiry and exploration capabilities, reflection and reasoning abilities, ethic and safety considerations, as well as practice proficiency. The SciKnowEval dataset includes 50K biology and chemistry-related problems along with their corresponding solutions.
@@
Dataset Last updated Subset #Item Domain Type Capability Language
MedQA-USMLE (Jin et al., 2020) 2021.04 General Medicine 61.0K Medical Multiple choice College Both
MedMCQA (Pal et al., 2022) 2022.03 General Medicine 193.1K Medical Multiple choice College English
JAMA Clinical (Chen et al., 2024a) 2024.03 General Medicine 1524 Medical Multiple choice College English
CMtMedQA (Yang et al., 2023b) 2023.08 General Medicine 68K Medical Multi-turn Dialogue College Chinese
huatuo-26M (Li et al., 2023f) 2023.03 General Medicine 26M Medical QA College Chinese
@@
Basic-chemistry 2909 Chemistry
@@
GAOKAO-Bench (Zhang et al., 2023c) 2023.11 Biology 266 Biology Multiple choice / QA Mixed Chinese
Chemistry 133 Chemistry
SciKnowEval (Feng et al., 2024) 2024.06 Biology 27730 Biology Multiple choice / QA / True or flase Mixed English
Chemistry 22250 Chemistry
Bioinfo-Bench-QA (Chen and Deng, 2023) 2023.10 - 150 Biology Multiple choice Post-college
BLURB (Gu et al., 2021) 2020.07 - 648k Biology Multiple NLP tasks Mixed English
PubMedQA (Jin et al., 2019b) 2019.09 - 273.2k Biology True or false College
SciBench (Wang et al., 2023c) 2023.07 - 272 Chemistry QA College